### v1

# Activated solids: Spontaneous deformations, non-affine fluctuations, softening, and failure

Abstract
We investigate spontaneous deformations in activated solids by analyzing crystal deformations through non-affine fluctuations. A scaling law for non-affinity is derived, showing it scales quadratically with activity and inversely with density. At high activity, the non-affine parameter diverges, signaling defect proliferation and melting to a hexatic phase. This softening is marked by reduced shear modulus, solid order, and hexatic order. At a higher activity, the hexatic melts to a fluid. We propose locally tuned activity to control non-affinity, demonstrated through numerical simulations. Our predictions are testable in active colloids performing persistent random motion.
I Introduction
The study of active solids is inspired by biological systems like cell assemblies, tissues, and bacterial biofilms, as well as artificial ones such as robotic solids [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. These systems are out of equilibrium at microscopic scales, with elements that locally consume and dissipate energy, generating propulsion or stress and breaking detailed balance. Unlike crystalline solids, which break continuous translational and rotational symmetries in equilibrium, amorphous solids can be stabilized by compression and are governed by local force and torque balance [16, 17, 18, 19]. In active solids maintaining such balance is essential for stability, as any departure can lead to deformation and flow.
Drawing inspiration from nature, various artificial micro- and nano-swimmer systems have been developed over last decades [20, 21]. These self-propelled colloids move by harnessing ambient energy – optical, thermal, electrical, or chemical – through an auto-catalytic process. Their behavior can be modeled by the active Brownian particle (ABP) framework, where particles move at a constant speed in their direction of motion, with random reorientations. This paper examines the impact of such activity on a sterically stabilized solid.
Equilibrium solids display low-energy phonon modes, described by displacement fields and elasticity theory. At higher excitations, topological defects, manifested as singularities in the displacement field, emerge. This framework not only captures the behavior of plastic deformation and yielding in various solids [22, 23, 24], but also offers a robust description of two-dimensional solid melting from quasi-long ranged ordered to disordered phase within the BKTHNY theory, driven by the proliferation of topological defects [25, 26, 27, 28, 29]. While the displacement field approach can be applied to amorphous solids with limited success [30], the absence of broken translational symmetry complicates distinguishing smooth from singular displacements [31, 32]. Instead, atomic configurations with localized, non-singular displacements – both proposed [33, 30] and observed experimentally [34] – are thought to function analogously to lattice defects in these materials.
A generalized approach to classifying atomic displacements in solids have been developed over the last decade [35, 36, 37, 38, 39, 40, 41, 42, 43]. Building on concepts used to identify local plastic events in structural glasses [17], it has been shown that atomic displacements within a coarse-graining region can be projected onto two orthogonal subspaces. The affine component represents homogeneous linear transformations, such as isotropic expansion, shear strains, and rotations, while the nonaffine component captures the residual displacements. These components are linearly independent [36]. Nonaffine modes have been linked to defect precursors in various lattice structures [42], bridging the projection formalism with conventional plasticity theories. This approach has also addressed fundamental questions in solid rigidity [41] and dynamic yield strength in nanoscale crystals [44]. Experimentally, nonaffine rearrangements have been observed to correlate with structural anisotropy in metallic glasses [45] and colloidal crystals [46].
In this paper, we explore spontaneous deformations in active solids using the concepts discussed above. The study of active matter [21, 20] has revealed novel non-equilibrium phenomena, such as flocking [47, 48], pattern formation and instabilities in active polar or apolar nematics [49, 50, 51, 52, 53], and motility-induced phase separation in active Brownian particles [54, 55, 56, 57], with their mechanisms understandable as low density phenomena [58, 59, 55, 60]. Experimental realizations involve, apart from natural flocks and swarms, rely on active colloids, active granular or robotic systems such as hexbugs [20, 61, 62, 63, 64, 65, 66, 67, 12, 13]. At high densities, active systems can exhibit liquid crystalline, jammed, or crystalline states [68, 60, 69, 70, 71, 72, 15]. Recent studies also highlight extreme spontaneous deformations in active crystals [73]. In this study, we characterize crystal deformations using non-affine fluctuations, deriving a scaling law in terms of activity $\Lambda$ and density. Our scaling argument predicts a $\Lambda^{2}$ dependence, confirmed over three decades through detailed numerical simulations – one of the key achievements of this paper. The non-affine parameter diverges at large $\Lambda$, signaling defect proliferation and activated melting of the solid to a hexatic phase, marked by vanishing shear modulus, solid order, and eventually the hexatic order. The softening and melting are further characterized analyzing the proliferation of topological defects. Finally, we propose locally tuned activity as a means to control non-affinity and the mechanical properties of a solid.
II Model
We consider an excluded volume solid in two dimensions characterized by a triangular lattice and quasi-long-range order, composed of $N=n_{x}n_{y}$ particles interacting via the Weeks-Chandler-Andersen (WCA) potential [74, 75], $U(r_{ij})=4\epsilon[{(\sigma/r_{ij})^{12}}-(\sigma/r_{ij})^{6}]+\epsilon$ for $|r_{ij}|<r_{c}=2^{1/6}\sigma$ and $U(r_{ij})=0$ otherwise. The system is maintained at high density $\rho=N/L_{x}L_{y}$ with mean inter-particle separation $a=[{2}/{\sqrt{3}\rho}]^{1/2}$. The equilibrium phase diagram, based on the WCA potential, reveals that for fixed temperature $k_{B}T/\varepsilon=1.0$, the solid undergoes a continuous BKTHNY transition to a hexatic phase at $\rho\sigma^{2}=0.92$, followed by first-order melting to a liquid at $\rho\sigma^{2}=0.906$ [76].
We investigate the behavior of an equilibrium solid at high density ($\rho\sigma^{2}\geq 1$) under self-propulsion, modeled by Active Brownian Particles (ABP). Each particle moves with self-propulsion speed $v_{0}$ in a direction $\hat{\mathbf{n}}_{i}={\cos\theta_{i}(t),\sin\theta_{i}(t)}$, with the angle $\theta_{i}(t)$ evolving via rotational diffusion. The dynamics of the $i$-th particle is governed by the following coupled Langevin equations:
$d{\bf r}_{i}(t)$ $=v_{0}\hat{\bf n}_{i}(t)dt-\mu\nabla_{i}\sum_{j\epsilon R_{i}}U(r_{ij})dt+\sqrt{2D_{t}dt}\,d{\bf B}_{i}(t)$
$d{\theta_{i}}(t)$ $=\sqrt{2D_{r}dt}\,dB^{r}_{i}(t),$ (1)
where $D_{t}$ and $D_{r}$ are the translational and rotational diffusivities, and $d\mathbf{B}_{i}(t)$ and $dB_{i}^{r}(t)$ are independent and univariate Gaussian white noise terms with mean zero. The Euler-Maruyama method is used to directly integrate the dimensionless forms of the above equations (see Appendix-A) in numerical simulations. The units of energy, length, and time are set by $\epsilon$, $\sigma$, $\tau_{u}=\sigma^{2}/\mu\epsilon$. The system property is controlled by the following dimensionless parameters: activity $\Lambda=v_{0}\sigma/\mu\epsilon$, translational diffusivity $\tilde{D}_{t}=D_{t}/\mu\epsilon$, and orientational diffusivity $\tilde{D}_{r}=D_{r}\sigma^{2}/\mu\epsilon$. In this paper, we focus on the impact of activity $\Lambda$, by setting $\tilde{D}_{r}=3$, $\tilde{D}_{t}=1$, and $\rho\sigma^{2}=1.1$, unless otherwise specified.
III Non-affine parameter
The non-affine fluctuation $\chi_{0}$ at lattice position ${\mathbf{R}_{0}}$ is defined within a coarse-graining volume $\Omega$ containing neighbors $\{\mathbf{R}_{i}\}$ (see Fig. 1c). Instantaneous positions ${\mathbf{r}_{0}}$ and $\{\mathbf{r}_{i}\}$ yield displacements $\mathbf{u}_{0}=\mathbf{r}_{0}-\mathbf{R}_{0}$ and $\mathbf{u}_{i}=\mathbf{r}_{i}-\mathbf{R}_{i}$ relative to reference coordinates $\{\mathbf{R}_{i}\}$ on an ideal triangular lattice. The local strain matrix ${\cal E}$ is obtained by minimizing the mean-square difference $D^{2}$ between the observed relative displacements $\mathbf{\Delta}_{i}=\mathbf{u}_{i}-\mathbf{u}_{0}$ and estimated displacements ${\cal E}[\mathbf{R}_{i}-\mathbf{R}_{0}]$ [17],
$\chi_{0}=\sum_{i}\left(\mathbf{\Delta}_{i}-{\cal E}[\mathbf{R}_{i}-\mathbf{R}_{0}]\right)^{2}.$ (2)
Note that $D$ has the dimension of length. After minimization, the $\alpha\beta$-th component of the affine strain is given by
${\cal E}^{m}_{\alpha\beta}=\sum_{\gamma}x_{\alpha\gamma}y_{\beta\gamma}^{-1}-\delta_{\alpha\beta},$ (3)
where
$x_{\alpha\beta}$ $=\sum_{i=1}^{N_{\Omega}}\left(r_{i\alpha}-r_{0\alpha}\right)\left(R_{i\beta}-R_{0\beta}\right),$
$y_{\alpha\beta}$ $=\sum_{i=1}^{N_{\Omega}}\left(R_{i\alpha}-R_{0\alpha}\right)\left(R_{i\beta}-R_{0\beta}\right).$
The residual $D^{2}$ value using the minimized ${\cal E}$ denoted by ${\cal E}^{m}$ in Eq.(3) gives the non-affine fluctuation $\chi_{0}$ of the 00-th particle [17, 35]:
$\chi_{0}=\sum_{i}\left(\mathbf{\Delta}_{i}-{\cal E}^{m}[\mathbf{R}_{i}-\mathbf{R}_{0}]\right)^{2}.$ (4)
A projection operator formalism [36] showed that this procedure projects the relative displacement field $\Delta_{i}$ onto mutually orthogonal subspaces of the affine strain tensor ${\cal E}^{m}$ and the non-affine parameter $\chi_{0}$. The global non-affine parameter $X=N^{-1}\sum_{i=1}^{N}\chi_{i}$ is the system average with $\chi_{i}$ denoting non-affine parameter associated with $i$-th particle. Note that $\chi$ and $X$ have dimensions of length squared, which we use in the next section to develop a scaling argument.
Figure 1: A schematic illustration of the activated solid. (a) An ideal triangular lattice that serves as the reference for calculation of non-affinity $\chi$. (b) A representative configuration at $\tilde{D}_{t}=1.0$ and $\Lambda=8.0$, is shown color coded by local non-affine parameter $\chi$; see the color bar. (c) A single coarse-graining region $\Omega$, which includes one central particle with reference position $\mathbf{R}_{0}$ (cyan) and actual location $\mathbf{r}_{0}$ (dark green). The reference (cyan, $\{\mathbf{R}_{i}\}$) and actual locations (dark green, $\{\mathbf{r}_{i}\}$) of its six nearest neighbors are shown as well.
III.1 Scaling analysis
In an equilibrium harmonic solid with spring constant $k$ and temperature $k_{B}T$, the characteristic length scale of displacement fluctuations is $\ell^{2}=\frac{k_{B}T}{k}$, and the mean non-affine parameter is $\langle\chi\rangle_{eq}\sim\ell^{2}$ [36]. With translational diffusion $D_{t}=\mu k_{B}T$, similar scaling is expected in our system, where the elastic constant $k$, derivable from $U^{\prime\prime}(a)$, increases with density. The bulk, shear, and Young’s moduli of a two-dimensional solid increase non-linearly with density [77], predicting an associated decrease in $\langle\chi\rangle$. At melting, the non-affine parameter is expected to diverge as the shear modulus ${\cal G}$ vanishes. Thus, the expression can be generalized to $\langle\chi\rangle_{eq}\sim\frac{D_{t}}{{\cal G}}$, with ${\cal G}$ increasing with density $\rho$ and vanishing at the melting point. Ignoring non-linearities for simplicity, the following emerges:
${\cal G}(\rho)$ $=$ $a(\rho-\rho_{m})\quad\text{for}\quad\rho>\rho_{m},$ (5)
[…]
with $\rho_{m}$ representing the melting density.
In this paper, we first consider the impact of activity on a solid maintained at a fixed temperature and density. As activity increases, the persistence length $\ell_{p}=v_{0}/D_{r}$ dominates system’s behavior. The effective diffusivity of ABPs in two dimensions is $D_{\rm eff}=D_{t}+v_{0}^{2}/2D_{r}$ [78, 79], which includes the contribution from activity. The non-affine parameter is expected to increase with activity as $\langle\chi\rangle\sim D_{\rm eff}/{\cal G}$ which in dimensionless form is:
$\langle\chi\rangle={\cal G}^{-1}(\tilde{D}_{t}+\Lambda^{2}/2).$ (6)
In active systems, the melting point $\rho_{m}$ depends on $\Lambda$ [15], affecting the density dependence of ${\cal G}$ in Eq.(5). Beyond that, the shear modulus decreases with activity as
${\cal G}={\cal G}_{0}-{\cal G}_{1}\Lambda^{2},$ (7)
as shown in a later section using numerical simulations. However, the $\Lambda$ dependence of ${\cal G}$ does not modify the dominant pre-melting scaling $\langle\chi\rangle\sim\Lambda^{2}$ appreciably.
In summary, for densities $\rho>\rho_{m}$ and activity $\Lambda<\Lambda_{c}$, the following scaling forms are expected:
$\langle\chi\rangle\sim(\rho-\rho_{m})^{-1}\quad\text{and}\quad\langle\chi\rangle\sim\Lambda^{2}.$ (8)
We compare these predictions with numerical simulations below.
Figure 2: (a) Evolution of the global non-affine parameter $X(t)$ at Λ=4, $\Lambda=4,\,10$. (b) Plot of $\langle\Delta X\rangle/a_{\rho}$ versus $\Lambda$ for $N=4096$ particles at varying densities $\rho$, where $\Delta X=X(\Lambda)-X_{eq}$, and $X_{eq}$ is the equilibrium value of global non-affinity. Filled markers represent the athermal system ($\tilde{D}_{t}=0$), and unfilled markers represent the thermal system ($\tilde{D}_{t}=1.0$). Data are fitted to $a_{\rho}\,\Lambda^{\nu}$ with best-fit exponents $\nu=2.00\pm 0.01$ (athermal) and $\nu=2.11\pm 0.06$ (thermal). Inset: density-dependent prefactor $a_{\rho}$ decreases with density as $(\rho-\rho_{0})^{-1}$, with $\rho_{0}\sigma^{2}=0.947\pm 0.008$ (athermal) and $\rho_{0}\sigma^{2}=0.971\pm 0.004$ (thermal).
Figure 3: (a): Shear modulus $\langle\mathcal{G}\rangle$ averaged over $n_{\mathrm{lin}}$ stress–strain curves showing linear response, plotted against $\Lambda$ for a system at density $\rho\sigma^{2}=1.1$. (b), (c), (d): Variation of solid order $\langle\psi_{\mathbf{G}}\rangle$, hexatic order $\langle\psi_{6}\rangle$, and defect fraction $\langle\delta_{d}\rangle$ with $\Lambda$; respective variances are shown in the insets. $\langle(\Delta\psi_{\mathbf{G}})^{2}\rangle$ peaks at $\Lambda=10.0$, $\langle(\Delta\psi_{6})^{2}\rangle$ shows a peak at $\Lambda=15.0$. Whereas $\langle(\Delta\delta_{d})^{2}\rangle$ exhibits two kinks at $\Lambda=10.2$ and $15.0$.
III.2 Scaling verification via numerical simulations
For this purpose, we analyze the global non-affine parameter $X=N^{-1}\sum_{i}^{N}\chi_{i}$. Its evolution reaches a steady state for $\Lambda<10$, but diverges for $\Lambda>10$ due to irreversible reorganization of particle positions. Examples of such trajectories of $X(t)$ at $\Lambda=4$ and $\Lambda=10$ are shown in Fig. 2(a). The steady-state regime is used to test the scaling hypothesis presented above.
We now consider the deviation of the global and steady-state non-affine parameter $\Delta X(\Lambda,\rho,T)=X(\Lambda,\rho,T)-X_{eq}(\rho,T)$ from equilibrium. Simulations with $N=4096$ ABPs, $\Lambda\in[0.5,8.0]$, $\rho\in[1.0,1.4]$, and $\tilde{D}_{t}=1.0$ yield a reasonable data collapse when data from different densities are plotted as $\langle\Delta X\rangle/a_{\rho}$ versus $\Lambda$. The excess non-affinity shows a scaling form $\langle\Delta X\rangle/a_{\rho}\sim\Lambda^{\nu}$ with $\nu=2.112\pm 0.006$; see Fig. 2(b). The inset in Fig. 2(b) shows the density dependence $a_{\rho}\sim(\rho-\rho_{0})^{-1}$ with $\rho_{0}\sigma^{2}=0.971\pm 0.004$, surpassing the equilibrium melting point $\rho^{eq}_{m}\sigma^{2}=0.92$ [76], reflecting active softening.
At $\Lambda\lesssim 1$, thermal fluctuations dominate over activity, making the $\Lambda$ dependence on the non-affine parameter difficult to discern. To address this, we set $\tilde{D}_{t}=0$ and explore the range $\Lambda\in[10^{-2},8]$. The results (filled symbols in Fig. 2) collapse onto a single curve across densities, resembling the $\tilde{D}_{t}=1.0$ case. Fitting the power law $a_{\rho}\,\Lambda^{\nu}$ gives $\nu=2.00\pm 0.01$, validating our predicted scaling for both thermal and athermal systems and confirming $\Delta X\sim\Lambda^{2}$ across three decades. The density-dependent prefactor $a_{\rho}\sim(\rho-\rho_{0})^{-1}$, with $\rho_{0}\sigma^{2}=0.947\pm 0.008$ (inset of Fig. 2(b) ). Though density scaling remains unchanged, $\rho_{0}\sigma^{2}$ is suppressed in the absence of thermal noise, as expected. The above analysis confirms both scaling forms presented in Eq. (8). The robustness of the scaling properties is further verified simulating larger system sizes in Appendix-B.
Figure 4: (a) Probability distribution $P(X)$ for $\rho=1.1$ at different $\Lambda$ values. Both mode and variance increase with activity. (b) $P(X)$ for different system sizes at $\Lambda=4$. Variance $\sigma^{2}$ of $P(X)$ showing $\sigma^{2}\sim N^{-1}$.
III.3 Probability distribution of $X$
In Figure 4, we explore the probability distributions of the global non-affine parameter $X$ for an ABP solid at $\tilde{D_{t}}=1$ and $\rho\sigma^{2}=1.1$. Figure 4a compares $P(X)$ for $N=4096$ particles across different $\Lambda$ values, showing that as $\Lambda$ increases, the mode of the distribution shifts to larger $X$ (consistent with the $\langle X\rangle\sim\Lambda^{2}$ scaling) and broadens with higher activity. This behavior capturing softening even in the absence of defect formation (as we show later, the onset of defect formation appears at $\Lambda\approx 10$). As shown in Fig. 4b for $\Lambda=4$ and $N$ ranging from 1024 to 16384, the probability distribution $P(X)$ is nearly Gaussian with variance $\sigma^{2}\sim 1/N$, consistent with the central limit theorem for the mean of uncorrelated entities. Similar scaling behavior for fluctuations in $X$ was observed before in equilibrium solids [39].
IV Analyzing structural indicators
Thus far, we have characterized the softening of a high-density solid through the non-affine parameter $X$ and its scaling with the Péclet number and density. In this section, we extend our analysis to include elastic and structural indicators, such as the shear modulus, solid order, and hexatic order, to further explore the active softening of solid. We perform this analysis using $\rho\sigma^{2}=1.1$ and $\tilde{D}_{t}=1.0$.
IV.1 Shear modulus
Within linear response, the shear modulus is given by $\mathcal{G}=\sigma^{tot}_{xy}/{\cal E}_{xy}$. We compute this from numerical simulations by applying a shear strain ${\cal E}_{xy}$ and measuring the change in total shear stress $\sigma_{xy}^{\mathrm{tot}}$ accounting for trajectories displaying linear response; see Appendix-C. The stress tensor has contribution from interaction calculated using the usual virial term and swim stress due to activity. The standard virial expression due to interaction gives,
$\sigma^{vir}_{\alpha\beta}=\frac{1}{dV}\sum_{i=1}^{N}\sum_{j=1}^{neb(i)}f_{ij,\alpha}\,r_{ij,\beta},$ (9)
where, $V$ denotes the system volume in $d=2$ dimensions, $f_{ij}$ is the interaction force between particles $i$ and $j$ with separation $r_{ij}$, and $\alpha$, $\beta$ denote the cartesian components of the vectors. On the other hand, the summation index $j$ runs over all interacting neighbors of the $i^{\mathrm{th}}$ particle. The active swim stress can be expressed as [80],
$\sigma^{sw}_{\alpha\beta}=\frac{\gamma v_{0}}{D_{r}V}\sum_{i=1}^{N}\big{\langle}\dot{r}_{i,\alpha}\,u_{i,\beta}\big{\rangle}.$ (10)
The total shear stress is then calculated as $\sigma^{tot}_{xy}=\sigma^{vir}_{xy}+\sigma^{sw}_{xy}$. Appendix-C outlines details of the shear modulus calculation.
Figure 3(a) shows the change in mean shear modulus, $\cal G$, with standard errors, as a function of activity $\Lambda$. We observe a steady decrease in it from the equilibrium value indicating softening of the solid with activity. As activity increases, fraction of trajectories following linear response decreases to vanish near $\Lambda=10$; see Appendix-C. The decrease in $G$ with activity can be described by ${\cal G}={\cal G}_{0}-{\cal G}_{1}\Lambda^{2}$ with ${\cal G}_{0}=109.84$ and ${\cal G}_{1}=0.094$. To further explore this softening and the fluidization transition, we compute structural indicators, including solid order $\psi_{\mathbf{G}}$, hexatic order $\psi_{6}$, and defect fraction $\delta_{d}$.
Figure 5: The structure factor $S(\mathbf{q})$ at $\mathrm{\Lambda}=4$ (a), $10$ (b), $10.2$ (c), $10.5$ (d), $15$ (e), and $20$ (f).
IV.2 Solid order
The structure factor of the system is given by
$S({\bf q})=\langle\rho_{\bf q}\rho_{-{\bf q}}\rangle$ (11)
where $\rho_{\bf q}=\frac{1}{N}\sum_{i=1}^{N}\exp(i{\bf q}\cdot{\bf r}_{i})$. Plots of $S({\bf q})$ at representative $\Lambda$ values are shown in Fig.5. In a perfect triangular lattice, quasi-Bragg peaks emerge at ${\bf q}_{p}=(0,\pm 2\pi/a_{y}),(\pm 2\pi/a,\pm\pi/a_{y})$, as shown in Fig. 5(a),(b). The solid order parameter $\langle\psi_{\bf G}\rangle$ averages $S({\bf q})$ over the six quasi-Bragg peaks at ${\bf G}:=\{{\bf q}_{p}\}$. Fig. 5(b) shows the variation of $\langle\psi_{\bf G}\rangle$ with $\Lambda$. The order parameter $\langle\psi_{\bf G}\rangle$ vanishes as the solid melts. At the melting point $\Lambda=10.0$ the variance of the order parameter $\langle(\Delta\psi_{\bf G})^{2}\rangle$ peaks (inset, Fig. 3(b) ). This analysis shows that the quadratic scaling of non-affinity with activity is observed across the entire activity range before solid melting.
As the active solid softens and melts with increasing $\Lambda$, the quasi-Bragg peaks flatten and merge (Fig. 5(c)-(f) ), similar to equilibrium 2D melting [76]. The $S({\bf q})$ in Fig. 5(c),(d) are characteristic of a hexatic. The six fold symmetry softens and starts to merge in Fig. 3(e) as the hexatic melts. Finally, at $\Lambda=20$, $S({\bf q})$ gets the uniform ring reflecting isotropy of the fluid; see Fig. 5(f). Therefore, $S({\bf q})$ suggests a two-stage melting from solid to hexatic to fluid.
Figure 6: Typical configurations with local hexatic order at $\Lambda=4$, $10.5$, and $15$: top row shows $|\psi^{k}_{6}|^{2}$ and bottom row shows $\mathrm{arg}(\psi^{k}_{6})$. Fig. 7 displays topological defects in the selected regions denoted by boxes in the figures.
IV.3 Hexatic order
This section examines the change in the hexatic order parameter to capture hexatic melting. Each particle in the system can be associated with a hexatic bond orientational order. For the $k$-th particle, it is defined as $\psi_{6}^{k}=(1/n)\sum_{i=1}^{n}e^{i6\theta_{kj}}$, where $\theta_{kj}$ is the bond angle between particle $k$ and its topological neighbor $j$ relative to the $x$-axis, and $n$ is the number of such neighbors [81]. In Fig. 3(c), we plot the mean amplitude of the hexatic order parameter of the system
$\langle\psi_{6}\rangle=\left\langle\left|\frac{1}{N}\sum_{k=1}^{N}\psi_{6}^{k}\right|^{2}\right\rangle,$ (12)
as a function of $\Lambda$. The hexatic order diminishes with $\Lambda$, but remains significant beyond $\Lambda=10$, suggesting the system stays in the hexatic phase past the solid melting point. At the hexatic melting point $\Lambda=15$, the hexatic order variance $\langle(\Delta\psi_{6})^{2}\rangle$ peaks sharply (inset of Fig. 3(c) ).
The local amplitudes $|\psi_{6}^{k}|^{2}$ and arguments arg($\psi_{6}^{k}$) of hexatic order for typical configurations across the transition are shown in Fig. 6 as the active solid softens and melts with increasing $\Lambda$. The drops in local hexatic amplitude, linked to hexatic orientation shifts, signal topological defect formation, as detailed in the following section.
Figure 7: (a) Voronoi tessellations for selected regions (see Fig. 6) at $\mathrm{\Lambda}=4.0$, $10.5$, and $15.0$. Particles are colored by their coordination number: green for five-fold, blue for seven-fold, and red for eight-fold, while sixfold-coordinated particles remain uncolored for clarity. (b) Regions containing defects are highlighted: (i) an isolated 5–7 defect; (ii) a pair of adjacent 5–7 defects; (iii) a cluster of three 5–7 defects surrounding a void; and (iv) a more complex defect cluster. (c) The fraction of sixfold-coordinated particles is plotted as a function of $\mathrm{\Lambda}$. (d) The fractions of particles with five- and seven-fold coordination are shown as functions of $\mathrm{\Lambda}$.
IV.4 Topological defects
We use Voronoi tessellations to identify each particle’s topological neighbors [81]. In a perfect lattice, each particle has 6 neighbors, with 5- or 7-fold deviations indicating topological defects. Minimal fluctuations in solid accommodate bound 5–7–5–7 dislocation pairs. The quilibrium BKTHNY theory links solid-hexatic melting to the unbinding of these pairs, creating dislocations, with hexatic melting driven by dislocation-unbinding to free 5- or 7-fold disclinations.
[…]
As activity $\Lambda$ exceeds $\Lambda=10.0$, the solid melting point, $\langle n_{6}\rangle$ decreases, while the defect fraction, $\langle n_{5,7}\rangle$, increases (see Fig. 7(c),(d)). The mean of total fraction of particles with non-six neighbors, $\delta_{d}=\frac{1}{N}\sum_{\nu\neq 6}N_{\nu}$, becomes non-zero at the solid melting point and increases with higher activity (Fig. 3(d)). The fluctuation in $\delta_{d}$ peaks twice: first at $\Lambda=10.2$, near the solid melting point, and again at $\Lambda=15.0$, at the hexatic melting point, reflecting enhanced defect fluctuations during both melting transitions (see inset of Fig. 3(d)).
V Local control of non-affinity and defect
Controlling local mechanical properties in solids has long been a challenge in physics and material science [82]. This can be achieved by manipulating non-affinity and defects, especially in colloidal crystals [83]. Previous studies in equilibrium solids used complex local laser fields [84] to drive non-affine parameters [39, 85]. Building on our study of non-affinity in activated melting, we propose a simpler method for local mechanical control: applying laser power to self-thermophoretic Janus colloids [64], inducing non-affinity and/or defects in an otherwise equilibrium solid. Localized activity control can also be achieved in other active colloidal or robotic systems [20, 86].
Figure 8: (a) Schematic of the simulation setup, where particles are selectively activated within a circular region of radius $r_{\rm beam}=30\sigma$ at the center of the simulation box, simulating laser irradiation. (b), (c) Snapshots of the system under $\Lambda=11$ at times $t=2\times 10^{6}\tau_{p}$ and $6.02\times 10^{6}\tau_{p}$, with particles colored by their local non-affinity $\chi_{i}$ and defects (particles with non-six neighbors) indicated. (d)–(f) Time evolution of the global non-affine deviation $\Delta X$ inside (blue) and outside (orange) the active region. (g)–(i) Evolution of defect fraction inside (blue) and outside (orange) the active region. Parameter values are $\Lambda=9$ (d,g), $\Lambda=10$ (e,h), and $\Lambda=11$ (f,i).
To study the above-mentioned effect near the solid–hexatic transition, we examine a collection of selectively activated particles coupled to a thermal heat bath with $\tilde{D}_{t}=1$, activated over a central circular region (Fig.8). Three activity levels are considered: below ($\Lambda=9$), at ($\Lambda=10$), and above ($\Lambda=11$) the transition. The system consists of $N=16384$ particles, initially arranged in a triangular lattice at density $\rho=1.1$ within a rectangular simulation box of size $131.14\sigma\times 113.57\sigma$. After sufficient equilibration, the central circular region of radius $r_{\rm beam}=30\sigma$ is selectively activated to mimic laser irradiation (Fig.8(a)). Particles are assigned activity $\Lambda_{i}(t)=\sigma_{i}(t)\Lambda$, with $\sigma_{i}=1$ inside the active zone and $\sigma_{i}=0$ outside. The value of $\sigma_{i}(t)$ is updated at each timestep based on particle positions.
A schematic of this setup is shown in Fig. 8(a), and representative snapshots during evolution for $\Lambda=11$ are shown in Fig. 8(b), (c), where particles are colored by their local non-affinity $\chi_{i}$. Pronounced softening and defect formation are observed inside the active region, while $\chi_{i}$ outside remains close to equilibrium. At early times, defects remain confined within the active zone; however, at later times, isolated defects begin to appear in the passive region as a result of migration and delayed annihilation of defects originating from the active area. To quantify global structural changes, we define $\Delta X=X(\Lambda)-X(0)$, the deviation of the global non-affine parameter from its equilibrium value, as defined in Sec. III.2. The evolution of this quantity in the active zone $\Delta X^{\rm(in)}$ and in the passive zone $\Delta X^{\rm(out)}$ is shown in Fig. 8(d)–(f). Below the transition ($\Lambda=9$), $\Delta X^{\rm(in)}$ fluctuates around a non-zero mean with a larger amplitude than in equilibrium, while $\Delta X^{\rm(out)}$ remains negligible. At transition ($\Lambda=10$), $\Delta X^{\rm(in)}$ diverges, while $\Delta X^{\rm(out)}$ stays flat, indicating localized melting. Above transition ($\Lambda=11$), this divergence occurs earlier followed by an increase in $\Delta X^{\rm(out)}$ as well, suggesting outward spread of non-affineness. The defect fraction evolution in Fig. 8(g)-(i) shows that for $\Lambda=9$, defects form and annihilate intermittently, while at and beyond the transition, defect fraction increases significantly inside and, to a lesser extent, outside the active region.
VI Discussion and Conclusion
In summary, this paper explored spontaneous deformations in active solids, characterizing crystal deformations through non-affine fluctuations and deriving a scaling law in terms of activity $\Lambda$ and density. Our scaling analysis and numerical data collapse predict that non-affinity scales as $\Lambda^{2}$ with activity and $(\rho-\rho_{m})^{-1}$ with density. At large $\Lambda$, the non-affine parameter diverges, signaling defect proliferation and the activated melting to a hexatic phase with vanishing shear modulus and order. This softening was further analyzed through topological defect proliferation. At higher activity, the hexatic melts to a fluid as captured by the change in structure factor and hexatic order parameter. Using the properties found from our studies, we finally proposed, locally tuned activity as a means to control non-affinity and mechanical properties. We numerically demonstrated the use of an isotropic activity patch in the solid. Such a patch locally softens the solid by increasing non-affinity, which reduces the shear modulus. Activating the solid in an anisotropic domain can modify elastic properties differently in each direction. Our predictions can be directly tested in active colloids. Additionally, locally controlling non-affinity via activity is simpler to implement in Janus colloid experiments than previous laser-controlled methods [39].
Author Contributions
The work began with a discussion between DC and SS, followed by initial non-affinity scaling simulations carried out by DD. PN set up and conducted all the final numerical simulations. DC supervised the project, and all authors contributed to writing the manuscript and discussed the results.
Data availability
The data are available from the authors upon reasonable request.
Acknowledgements.
DC acknowledges research grants from the Department of Atomic Energy (OM no. 1603/2/2020/IoP/R&D-II/15028), an Associateship at ICTS-TIFR, Bangalore, and expresses gratitude to MPIPKS, Dresden for hospitality during a two-month visit in 2024, where part of the research was carried out. The simulations were performed using resources at IOP Bhubaneswar. DC, DD, and PN sincerely thank Surajit Sengupta [87], who tragically passed away, for his advocacy of using non-affine fluctuations to characterize active solids, and dedicate this paper to his memory.
Appendix A Dimensionless equations
To perform numerical simulations, we use the dimensionless equations by applying the following substitutions: ${\bf r}_{i}\to\tilde{\bf r}_{i}={\bf r}_{i}/\sigma$, $t\to\tilde{t}=t/\tau_{u}=t\mu\epsilon/\sigma^{2}$, the dimensionless activity $\Lambda=v_{0}\tau_{u}/\sigma=v_{0}\sigma/\mu\epsilon$, the dimensionless translational diffusivity $D_{t}\to\tilde{D}_{t}=D_{t}\tau_{u}/\sigma^{2}=D_{t}/\mu\epsilon$, $\nabla_{i}\to\tilde{\nabla}_{i}=\sigma\nabla_{i}$, $U(r_{ij})\to\tilde{U}=U/\epsilon$, $\tilde{\mu}=\mu\epsilon\tau_{u}/\sigma^{2}=1$ and the dimensionless rotational diffusivity $\tilde{D}_{r}=D_{r}\sigma^{2}/\mu\epsilon$. By using these substitutions, one obtains the dimensionless form of the equations
$d\tilde{\bf r}_{i}(\tilde{t})$ $=\Lambda\,\hat{\bf n}_{i}(\tilde{t})d\tilde{t}-\tilde{\nabla}_{i}\sum_{j\epsilon R_{i}}\tilde{U}(r_{ij})d\tilde{t}+\sqrt{2\tilde{D}_{t}d\tilde{t}}\,d{\bf B}_{i}(\tilde{t})$
$d{\theta_{i}}(\tilde{t})$ $=(2\,\tilde{D}_{r}d\tilde{t})^{1/2}\,dB^{r}_{i}(\tilde{t}).$ (13)
We perform numerical simulations with $\tilde{D}_{r}=3$, $\tilde{D}_{t}=1$ for thermal, and $\tilde{D}_{t}=0$ for athermal cases.
Appendix B Scaling ansatz for $X$ at different system sizes
Figure 9: As in Fig. 2, we plot $\langle\Delta X\rangle$ for two system sizes ($N=4096,16384$) at density $\rho\sigma^{2}=1.1$, demonstrating that the scaling $\langle\Delta X\rangle\sim\Lambda^{2}$ holds for both the thermal ($\tilde{D}_{t}=1$) and athermal ($\tilde{D}_{t}=0$) cases at the larger system size.
In Fig. 2 of the main text, we showed the scaling behavior $\langle\Delta X\rangle\sim\Lambda^{2}$ for a system of size $N=4096$. To verify the robustness of this scaling, we present additional data for a larger system size ($N=16384$) at density $\rho\sigma^{2}=1.1$ in Fig. 9. We confirm that the quadratic scaling $\langle\Delta X\rangle\sim\Lambda^{2}$ remains valid for both thermal ($\tilde{D}_{t}=1$) and athermal ($\tilde{D}_{t}=0$) cases, consistently across system sizes.
Appendix C Obtaining shear modulus
Figure 10: (a) Shear stress, $\sigma_{xy}^{tot}$, vs. shear strain, $\mathcal{E}_{xy}=\gamma$, at $\Lambda=7$ for 24 ensembles without significant non-linearity or hysteresis. Different colors represent individual ensembles, with the black line showing the linear fit, where the slope gives the shear modulus, $\langle{\cal G}\rangle=105.7\pm 0.679$ in units of $\epsilon/\sigma^{2}$. (b) Linear response trajectories and (c) non-linear response trajectories exhibiting significant hysteresis and nonlinearity. Non-linear trajectories are excluded from shear modulus calculation. Raw data are shown as semi-transparent circular markers, with opaque square markers representing averages over 100 consecutive values of $\gamma$ and $\sigma^{\mathrm{tot}}_{xy}$. Markers are color-coded by shear direction: forward shear $0\rightarrow\gamma_{m}$ in red, backward shear $\gamma_{m}\rightarrow-\gamma_{m}$ in green, and $-\gamma_{m}\rightarrow 0$ in blue. (d) Fraction of linear ensembles, $n_{\mathrm{lin}}/n_{\mathrm{tot}}$, as a function of activity $\Lambda$.
For each $\Lambda$, we initialize with a triangular lattice and run simulations for $t_{\max}=10^{7}$ with a time step $dt=10^{-5}\tau_{p}$ until steady state is reached. We then apply three deformation stages:
[…]
at a strain rate ˙$\dot{{\cal E}}=10^{-2}D_{r}$ and ${\cal E}_{xy}^{\max}=0.035$. The shear modulus is a linear response coefficient; thus the chosen ${\cal E}_{xy}^{\max}$ is to ensure the system response stays within the linear regime for accurate shear modulus extraction.
For $\Lambda\leq 6$, all four initial configurations exhibit linear stress-strain behavior. For $\Lambda>6$, non-linear deviations appear. We generate $n_{\mathrm{tot}}\approx 100$ initial configurations, discarding those with nonlinearity for shear modulus analysis, which becomes more common at higher $\Lambda$. The shear modulus is then computed by averaging the linear response data from the remaining $n_{\mathrm{lin}}$ hysteresis-free ensembles, as shown in Fig. 10(a). Examples of linear and non-linear ensembles are shown in Figs. 10(b) and (c).
The fraction $n_{\mathrm{lin}}/n_{\mathrm{tot}}$ decreases with $\Lambda$, reaching zero at the melting point, $\Lambda=10.0$ (Fig. 10(d)). This decline for $\Lambda>6.0$ indicates activity-driven softening, marked by a departure from linear elasticity. Near $\Lambda=10.0$, both $n_{\mathrm{lin}}$ and the shear modulus approach zero, signaling a fluidization transition with the complete loss of linear response.